Kesken:

Nimi  Op.kirjan nro Esitys Kalvot Tehtävät Harj. työ AS
Teemu Karevaara  44885J  ok  ok
Jouni Seppänen  55936A  ok+ ok+  hyv.
Jukka Laurila  47029E  ok-  ok
Janne Sinkkonen  ?  ok+  ok
Jarkko Salojärvi  40734V  ok+ ok+
Olli-Pekka Rinta-Koski  36336K  ok+  ok
Ismo Haanaho  44143H  ok+  ok
Jarkko Venna  34110F  ok+  ok
Jörkki Hyvönen  44403A  ok  ok
Satu Nahkuri  ?  ok+  ok
Petri Loukasmäki  ?  ok+  ok

Suorittaneet 2.3.2001:

Nimi  Op.kirjan nro Esitys Kalvot Tehtävät Harj. työ AS
Ville Viitaniemi  44374L  ok  ok  k.hyv.  k.hyv. k.hyv.
Raul Lozano  55811K  ok-  ok  hyv.  hyv.  hyv.
Anne Patrikainen  50135R  ok+ ok+  k.hyv.  k.hyv. k.hyv.
Juha Alaluusua  46532B  ok+  ok  hyv.  hyv.  hyv.
Eero Vaarnas  46674M  ok+  ok  hyv.  hyv.  hyv.
Markku Roiha  29883T  ok+  ok  hyv.  hyv.  hyv.
Antonio Villena  55809H  ok+  ok  hyv.  hyv.  hyv.
Jukka Lindström  47798R  ok-  ok  hyv.  hyv.  hyv.
Mika Inki  47011F  ok  ok  k.hyv.  k.hyv. k.hyv.
Matti Aksela  44009F  ok  ok  hyv.  k.hyv.  hyv.
Harri Laakso  46589B  ok+  ok  hyv.  hyv.  hyv.
Julian Cobos  55810J  ok  ok  hyv.  hyv.  hyv.
Ville Könönen  94196R  ok  ok  k.hyv.  k.hyv. k.hyv.
Janne Nikkilä  41903K  k.hyv

Suorittaneet 9.1.2002 mennessä:

Nimi  Op.kirjan nro Esitys Kalvot Tehtävät Harj. työ AS
Vesa Siivola  43036B  ok  ok  hyv  k.hyv.